# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 6 Conclusion
Table S4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 (full data), along with the computation time (CT) in seconds for one VB iteration. The table also presents the DIC values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
Table S5 presents the posterior means and 95% credible intervals for selected model parameters obtained from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models applied to the simulated dataset 2, along with the computational cost for one VB iteration. The table also includes the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model. Corresponding posterior density plots of the parameters are provided in Section S7.2.1. Unlike the simulated dataset 1, which is skewed and heavy tails, the simulated dataset 2 is more symmetrical with lighter tails. The posterior mean estimates for parameters $\beta_{0}$, $\beta_{1}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ across all four models closely match their true values. Likewise, the posterior mean estimate of $\nu$ for SEM-t and YJ-SEM-t is close to the true value of $30$, while the posterior mean estimate of $\gamma$ for YJ-SEM-Gau and YJ-SEM-t is close to the true value of $1$. The overall accuracy of parameter estimation across all models is further supported by the nearly identical $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, with SEM-Gau having the lowest values, indicating the best fit for the simulated dataset 2, as expected.
According to the computation times presented in Tables S4 and S5, VB algorithms for both SEM-t and YJ-SEM-t, which incorporate Student’s $t$ errors, require significantly more computation time compared to SEM-Gau and YJ-SEM-Gau. This is because, in models with $t$ errors, the VB optimisation is substantially more complex than in models with Gaussian errors. Specifically, for models with Student’s $t$ errors, the set of model parameters is $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, whereas for models with Gaussian errors, it simplifies to $\boldsymbol{\xi}=\boldsymbol{\phi}$. Here, $\boldsymbol{\tau}$ is a latent variable vector of length $n$; see Section 2.1 of the main paper.
@@
Table S5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 (full data), along with the computation time (CT) in seconds for one VB iteration. The table also presents the DIC values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
S7 Additional figures for simulation study in Section 4
This section provides additional figures related to the simulation study presented in Section 4 of the main paper, and Section S6.
S7.1 Comparison of VB and HVB methods
Section 4.1 of the main paper compares the estimated posterior densities of the model parameters for the YJ-SEM-Gau model using VB (including HVB) and HMC (Neal,, 1996) methods, both with and without missing data, based on a simulated dataset. To generate the data, the model parameters are set as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from discrete uniform values between $-3$ and $3$ (excluding 00). The error variance parameter is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autoregressive parameter is $\rho=0.8$. The Yeo-Johnson transformation parameter is set to $\gamma=1.25$, which introduces moderate left skewness in the simulated response variable. The kernel density plot of the simulated response variable is provided in Figure S1. As expected, the density is moderately skewed to the left.
Figure S1: Kernel density plot of the simulated response variable generated using the YJ-SEM-Gau model with $\gamma=1.25$ in the simulation study of Section 4.1 of the main paper.
S7.2 Evaluating the performance of proposed SEMs
S7.2.1 Simulation study with full data
This section presents further figures from Section S6.
@@
Figures S2 and S3 compare the posterior densities of selected model parameters obtained using the proposed VB method, applied to different SEMs (without missing values) for simulated dataset 1 and simulated dataset 2, respectively.
As shown in Figure S2, the YJ-SEM-t model is able to recover the true parameter values more accurately than the other three models, as expected, since the simulated dataset 1 was simulated using the YJ-SEM-t model. In contrast, as shown in Figure S3, all four models successfully recover the true values of the parameters.
Figure S2: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 1, using the VB method. The vertical dotted lines indicate the true parameter values.
Figure S3: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 2, using the VB method. The vertical dotted lines indicate the true parameter values.
S7.2.2 Simulation study with missing data
This section presents further figures from Section 4.2 of the main paper.
Figures S4 and S5 compare the posterior densities of selected model parameters estimated using the proposed HVB method, applied to different SEMs with missing values, for simulated dataset 1 and simulated dataset 2, respectively.
As shown in Figure S4, the YJ-SEM-t model is able to recover the true parameter values more accurately than the other three models, as expected, since the simulated dataset 1 was simulated using the YJ-SEM-t model. In contrast, as shown in Figure S5, all four models successfully recover the true values of the parameters.
@@
Figure S4: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 1 with missing data, using the HVB method. The vertical dotted lines indicate the true parameter values.
@@
Figure S5: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 2 with missing data, using the HVB method. The vertical dotted lines indicate the true parameter values.
The left panel of Figure S6 shows the kernel densities of the true missing values ($\mathbf{y}_{u}$) and the kernel density of posterior means of the missing values at locations $s_{1},s_{2},\ldots,s_{n}$, estimated from different SEMs for simulated dataset 2. All four models produce nearly identical densities of posterior means of missing values that closely match the density of the true missing values.
The right panel of Figure S6 presents the posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by each SEM. The true value of $\textbf{y}_{u_{\textrm{max}}}$ lies within the posterior distribution estimated by all models. These results are consistent with expectations, as all four models produce similar parameter estimates and DIC values (see Table 5).
@@
Figure S6: Left panel: The kernel densities of the true missing values ($\textbf{y}_{u}$) and the posterior means of the missing values ($\textbf{y}_{u}$) obtained using the HVB-AllB method for different SEMs for the simulated dataset 2 with missing values.
Right panel: The posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by different SEMs for the simulated dataset 2 with missing values. The true value of the maximum missing value is indicated by the vertical line.
@@
Figure S7 compares the posterior means and standard deviations of the missing values ($y_{u_{1}},\ldots,y_{u_{n_{u}}}$) estimated by different SEMs for the simulated dataset 1 with missing values. The left panel compares the posterior means of the predicted missing values with their true values. The right panel compares the posterior standard deviations of the missing values estimated by the YJ-SEM-t model (x-axis) against those obtained from SEM-Gau, SEM-t, and YJ-SEM-Gau models (y-axis).
The posterior means of the missing values estimated by the YJ-SEM-Gau and YJ-SEM-t models are closer to the true missing values compared to those estimated by the SEM-Gau and SEM-t models. Additionally, the posterior standard deviations from the YJ-SEM-Gau and YJ-SEM-t models are nearly identical, whereas those from the SEM-Gau and SEM-t models deviate from the standard deviations obtained by the YJ-based models.
@@
Figure S7: Comparison of posterior means and standard deviations of missing values for the simulated dataset 1: The left plot shows the comparison of the posterior means of predicted missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values. The right plot compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
@@
Figure S8 compares the posterior means and standard deviations of the missing values ($y_{u_{1}},\ldots,y_{u_{n_{u}}}$) estimated by different SEMs for the simulated dataset 2 with missing values. The left panel compares the posterior means of the predicted missing values with their true values. The right panel compares the posterior standard deviations of the missing values estimated by the YJ-SEM-t model (x-axis) against those obtained from SEM-Gau, SEM-t, and YJ-SEM-Gau models (y-axis).
The posterior means of the missing values estimated by all four models are close to the true values. Additionally, the posterior standard deviations from the SEM-Gau, SEM-t, and YJ-SEM-Gau models closely align with those from the YJ-SEM-t model, with YJ-SEM-Gau showing the closest match.
@@
Figure S8: Comparison of posterior means and standard deviations of predicted missing values for the simulated dataset 2: The left plot shows the comparison of the posterior means of missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values. The right plot compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
S8 Additional figures for real data application in Section 5
@@
Figures S10 and S11 compare the posterior densities of selected model parameters, estimated using the proposed VB and HVB-AllB methods, across different SEMs (without and with missing values) applied to the Lucas-1998-HP dataset.
Figure S10: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the Lucas-1998-HP dataset without missing values, using the VB method.
Figure S11: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the Lucas-1998-HP dataset with missing values, using the HVB-AllB method.
S9 Convergence analysis
@@
Section S9.1 presents the convergence analysis plots for the simulation study described in Section 4 of the main paper. Similarly, Section S9.2 provides the convergence analysis plots for the real-data example discussed in Section 5 of the main paper.
S9.1 Convergence analysis of simulation study
S9.1.1 Convergence analysis for Section 4.1
This section presents convergence plots from the simulation study designed to compare the performance of the VB approximation with the HMC algorithm, as described in Section 4.1 of the main paper. In this study, data are simulated from the YJ-SEM-Gau model, and the model is estimated using both VB and HMC methods.
Figure S12 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau model with full data, after excluding burn-in samples. Figure S13 displays the trajectories of the variational means for selected parameters across VB iterations. Both plots indicate that the algorithms have converged.
Figure S12: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau model (without missing values), after excluding burn-in samples.
Figure S13: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau model (without missing values) across VB iterations.
Next, missing values are introduced into the simulated dataset, and the YJ-SEM-Gau model is fitted using both the HVB-NoB and HMC algorithms.
Figure S14 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau model with missing values, after excluding burn-in samples. Figure S15 displays the trajectories of the variational means for selected parameters across HVB-NoB iterations. Both plots indicate that the algorithms have converged.
Figure S14: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau model (with missing values), after excluding burn-in samples.
Figure S15: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau model (with missing values) across HVB-NoB iterations.
S9.1.2 Convergence analysis for Section 4.2
This section presents convergence plots from the simulation study conducted to evaluate the performance of the proposed SEMs, as described in Section 4.2 of the main paper.
@@
S9.2 Convergence analysis of real example
Figure S20 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset (full data). It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the HVB-AllB algorithms for all models have successfully converged.
Figure S20: Trajectories of the variational means of selected parameters across HVB-AllB iterations for different SEMs fitted on the Lucas-1998-HP dataset without missing values.
Figure S21 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset, with missing values. It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the VB algorithms for all models have successfully converged.

### v2

@@ 9 Conclusion
We begin by comparing the estimated values of the fixed effects (specifically $\beta_{0}$ and $\beta_{1}$), the variance parameter $\sigma_{\textbf{e}}^{2}$, and the spatial autocorrelation parameter $\rho$, as these parameters are present in all four models. The posterior means for $\beta_{0}$, $\beta_{1}$ and $\rho$ are nearly identical for YJ-SEM-Gau and YJ-SEM-t, closely matching the true values of 1-1 for $\beta_{0}$, $1$ for $\beta_{1}$, and $0.8$ for $\rho$. In contrast, the estimates from SEM-Gau and SEM-t show significant deviations from the true values. The estimated posterior means of $\sigma_{\textbf{e}}^{2}$ from SEM-Gau, SEM-t, and YJ-SEM-Gau deviate considerably from the true value, while the posterior mean obtained from YJ-SEM-t is much closer to the true value. The models incorporating the YJ transformation, YJ-SEM-Gau and YJ-SEM-t, successfully recover the parameter $\gamma$. The degrees of freedom ($\nu$) estimated by the SEM-t and YJ-SEM-t models show slight deviations from the true value.
For the simulated dataset 1, the model fit assessment indicates that YJ-SEM-t is the best-performing model, as it achieves the lowest $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values among all four models. This outcome is expected, given that the dataset is simulated from YJ-SEM-t. The second-best model is YJ-SEM-Gau, which has the next lowest DIC values. Notably, the difference in DIC between YJ-SEM-t and YJ-SEM-Gau is significantly small, suggesting that YJ-SEM-Gau provides a model fit nearly as good as YJ-SEM-t for the dataset 1.
Figure S2 compares the posterior densities of selected model parameters obtained using the proposed VB method, applied to different SEMs for the simulated dataset 1. As expected, the YJ-SEM-t recovers the true parameter values more accurately than the other three models, since dataset 1 is simulated using the YJ-SEM-t.
Figure S2: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the simulated dataset 1 (without missing data), using the VB method. The vertical dotted lines indicate the true parameter values.
S8.2 Results for the dataset 2
Table S5 presents the posterior means and 95% credible intervals for selected model parameters obtained from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t applied to the simulated dataset 2, along with the computation cost for one VB iteration. The table also includes the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model. Unlike the simulated dataset 1, which is skewed and has heavy tails, the simulated dataset 2 is more symmetrical with lighter tails. The posterior mean estimates for parameters $\beta_{0}$, $\beta_{1}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ across all four models closely match their true values. Likewise, the posterior mean estimate of $\nu$ for SEM-t and YJ-SEM-t is close to the true value of $30$, while the posterior mean estimate of $\gamma$ for YJ-SEM-Gau and YJ-SEM-t is close to the true value of $1$. The overall accuracy of parameter estimation across all models is further supported by the nearly identical $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, with SEM-Gau having the lowest values, indicating the best fit for the simulated dataset 2.
Table S5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 (full data), along with the computation time (CT) in seconds for one VB iteration. The table also presents the DIC values for each model. Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
@@
Figure S3 compares the posterior densities of selected model parameters obtained using the proposed VB method, applied to different SEMs for simulated dataset 2. It can be seen that, all four models successfully recover the true values of the parameters.
@@
Figure S3: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the simulated dataset 2 (without missing data), using the VB method. The vertical dotted lines indicate the true parameter values.
S8.3 Summary and computation time comparison
According to the computation times presented in Tables S4 and S5, VB algorithms for both SEM-t and YJ-SEM-t, which incorporate Student’s $t$ errors, require significantly more computation time per iteration compared to SEM-Gau and YJ-SEM-Gau. This is because, in models with Student’s $t$ errors, the VB optimisation is substantially more complex than in models with Gaussian errors. Specifically, for models with Student’s $t$ errors, the set of model parameters is $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, whereas for models with Gaussian errors, it simplifies to $\boldsymbol{\xi}=\boldsymbol{\phi}$. Here, $\boldsymbol{\tau}$ is a latent variable vector of length $n$; see Section 2.1 of the main paper. In addition to the higher per-iteration cost, SEM-t and YJ-SEM-t also require more VB iterations to achieve convergence. Together, these factors lead to substantially higher computational costs for models with Student’s $t$ errors than for their Gaussian-error counterparts.
In terms of estimation performance, the results indicate that both YJ-SEM-Gau and YJ-SEM-t are particularly effective at capturing the key characteristics of non-Gaussian data. For simulated dataset 1, which exhibits skewness and heavy tails, YJ-SEM-t achieves the lowest DIC values, indicating the best fit, while YJ-SEM-Gau attains the second-lowest values. The differences between the DIC values of YJ-SEM-t and YJ-SEM-Gau are small, suggesting that YJ-SEM-Gau provides a model fit comparable to that of YJ-SEM-t. For simulated dataset 2, which is closer to Gaussian, all four models recover the true parameter values accurately, and the DIC values are nearly identical, with SEM-Gau achieving the lowest values and therefore providing the best fit.
S9 Additional figures for simulation study: Assessing the accuracy and robustness of proposed SEMs (missing data case in Section 7.1)
This section provides additional figures supporting the analysis in Section 7.1 of the main paper, which evaluates the accuracy and robustness of the proposed SEMs and HVB (HVB-AllB) methods using two simulated datasets with missing values.
S9.1 Comparison of posterior densities of model parameters
Figures S4 and S5 present posterior densities of selected model parameters from different SEMs, estimated using the proposed HVB-AllB method on simulated datasets 1 and 2 with missing values. Figure S4 shows that YJ-SEM-t recovers the true parameter values most accurately for simulated dataset 1, as expected, since this dataset was generated from the YJ-SEM-t. In contrast, Figure S5 demonstrates that, for simulated dataset 2, all four models recover the true parameter values well.
@@
Figure S4: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the simulated dataset 1 with missing data, using the HVB-AllB method. The vertical dotted lines indicate the true parameter values.
@@
Figure S5: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the simulated dataset 2 with missing data, using the HVB-AllB method. The vertical dotted lines indicate the true parameter values.
S9.2 Comparison of posterior densities of missing values
@@
Figure S6 compares the posterior means and standard deviations of the missing values $\textbf{y}_{u}={(y_{u_{1}},\ldots,y_{u_{n_{u}}})}^{\top}$ estimated by the HVB-AllB algorithms under different SEMs for the simulated dataset 1 with missing values. The left panel compares the posterior means of the predicted missing values with their true values. The right panel compares the posterior standard deviations of the missing values estimated under the YJ-SEM-t (x-axis) with those obtained from the SEM-Gau, SEM-t, and YJ-SEM-Gau (y-axis). The posterior means of the missing values estimated under the YJ-SEM-Gau and YJ-SEM-t are closer to the true missing values compared to those estimated under the SEM-Gau and SEM-t. The posterior standard deviations of the missing values from the YJ-SEM-Gau and YJ-SEM-t are nearly identical, whereas those from the SEM-Gau and SEM-t deviate from the standard deviations obtained by the models with YJ transformation.
@@
Figure S6: Comparison of posterior means and standard deviations of missing values for the simulated dataset 1: The left panel displays the comparison of the posterior means of predicted missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values. The right panel compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
@@
Figure S7 compares the posterior means and standard deviations of the missing values $\textbf{y}_{u}={(y_{u_{1}},\ldots,y_{u_{n_{u}}})}^{\top}$ estimated by the HVB-AllB algorithms under different SEMs for the simulated dataset 2 with missing values. The left panel compares the posterior means of the predicted missing values with their true values. The right panel compares the posterior standard deviations of the missing values estimated under the YJ-SEM-t (x-axis) with those obtained from the SEM-Gau, SEM-t, and YJ-SEM-Gau (y-axis). The posterior means of the missing values estimated under all four models are close to the true values. Additionally, the posterior standard deviations from the SEM-Gau, SEM-t, and YJ-SEM-Gau closely align with those from the YJ-SEM-t, with YJ-SEM-Gau showing the closest match.
@@
Figure S7: Comparison of posterior means and standard deviations of predicted missing values for the simulated dataset 2: The left panel displays the comparison of the posterior means of missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values. The right panel compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
The left four panels of Figure S8 compare the kernel density of the true missing values ($\textbf{y}_{u}$) with the kernel densities of the posterior means of the missing values estimated under different SEMs for the simulated dataset 2. All four models produce nearly identical densities of posterior means of missing values that closely match the density of the true missing values. The right panel of Figure S8 presents the posterior density of the maximum missing value, $\textrm{max}(\textbf{y}_{u})$, estimated under each SEM. The true value of $\textrm{max}(\textbf{y}_{u})$ lies within the posterior distribution estimated by all models. These results are consistent with expectations, as all four models produce similar parameter estimates and $\text{DIC}_{5}$ values (see Table 5 in Section 7.1 of the main paper). A corresponding figure for simulated dataset 1 is presented in Figure 4 in Section 7.1 of the main paper.
Figure S8: Left panel: The kernel densities of the true missing values ($\textbf{y}_{u}$) and the posterior means of the missing values ($\textbf{y}_{u}$) obtained using the HVB-AllB method for different SEMs for the simulated dataset 2 with missing values. Right panel: The posterior density of the maximum missing value, $\textrm{max}(\textbf{y}_{u})$, predicted by different SEMs for the simulated dataset 2 with missing values. The true value of the maximum missing value is indicated by the vertical line.
S10 Additional figures for real data application in Section 8
@@
Figures S10 and S11 compare the posterior densities of selected model parameters, estimated using the proposed VB and HVB-AllB methods, across the different SEMs applied to the Lucas-1998-HP dataset, without and with missing values, respectively.
Figure S10: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the Lucas-1998-HP dataset without missing values, using the VB method.
Figure S11: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the Lucas-1998-HP dataset with missing values, using the HVB-AllB method.
S11 Convergence analysis
@@
Section S11.1 presents convergence analysis plots for the simulation study described in Section 6 of the main paper. Section S11.2 provides the corresponding convergence analysis plots for the simulation study in Section 7 of the main paper (with missing data), as well as for the full-data case in Section S8. Finally, Section S11.3 reports the convergence analysis plots for the real-data example discussed in Section 8 of the main paper.
S11.1 Convergence analysis for simulation study: Assessing the accuracy of VB methods
This section presents convergence plots from the simulation study designed to compare the performance of the VB approximation with the HMC algorithm, as described in Section 6 of the main paper. In this study, data are simulated from the YJ-SEM-Gau, and the model is estimated using both VB and HMC methods.
Figure S12 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau with full data, after excluding burn-in samples. Figure S13 displays the trajectories of the variational means for selected parameters across VB iterations. Both plots indicate that the algorithms have converged.
Figure S12: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau (without missing values), after excluding burn-in samples.
Figure S13: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau (without missing values) across VB iterations.
Next, missing values are introduced into the simulated dataset, and the YJ-SEM-Gau is fitted using both the HVB-NoB and HMC algorithms.
Figure S14 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau with missing values, after excluding burn-in samples. Figure S15 displays the trajectories of the variational means for selected parameters across HVB-NoB iterations. Both plots indicate that the algorithms have converged.
Figure S14: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau (with missing values), after excluding burn-in samples.
Figure S15: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau (with missing values) across HVB-NoB iterations.
S11.2 Convergence analysis for simulation study: Assessing the accuracy and robustness of proposed SEMs
This section presents the convergence plots for the simulation study in Section 7 of the main paper (with missing data), as well as for the full-data case in Section S8.
@@
S11.3 Convergence analysis of real data application
This section presents the convergence analysis plots for the real data application discussed in Section 8 of the main paper.
Figure S20 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset (full data). It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the VB algorithms for all models have successfully converged.
Figure S20: Trajectories of the variational means of selected parameters across VB iterations for different SEMs fitted on the Lucas-1998-HP dataset without missing values.
Figure S21 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset, with missing values. It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the HVB-AllB algorithms for all models have successfully converged.